# Do Words Reflect Beliefs? Evaluating Belief Depth in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly shaping political discourse, yet their responses often display inconsistency when subjected to scrutiny. While prior research has primarily categorized LLM outputs as left- or right-leaning to assess their political stances, a critical question remains: Do these responses reflect genuine internal beliefs or merely surface-level alignment with training data? To address this, we propose a novel framework for evaluating belief depth by analyzing (1) argumentative consistency and (2) uncertainty quantification. We evaluate 12 LLMs on 19 economic policies from the Political Compass Test, challenging their belief stability with both supportive and opposing arguments. Our analysis reveals that LLMs exhibit topic-specific belief stability rather than a uniform ideological stance. Notably, up to 95% of left-leaning models’ responses and 89% of right-leaning models’ responses remain consistent under the challenge, enabling semantic entropy to achieve high accuracy (AUROC=0.78), effectively distinguishing between surface-level alignment from genuine belief. These findings call into question the assumption that LLMs maintain stable, human-like political ideologies, emphasizing the importance of conducting topic-specific reliability assessments for real-world applications.
@@ 1 Introduction
The growing influence of Large Language Models (LLMs) in public discourse has intensified scrutiny of how they encode and express political biases inherited from training data (Feng et al., 2023; Santurkar et al., 2023). While studies consistently demonstrate that modern LLMs exhibit left-libertarian leanings (Yang et al., 2024; Motoki et al., 2024), their responses to political topics remain alarmingly inconsistent (Ceron et al., 2024). For instance, models may advocate progressive taxation in one context yet contradict themselves when presented with counterarguments framed as authoritative claims (Chen et al., 2024; Anagnostidis & Bulian, 2024). This irrational compliance contrasts sharply with human ideological behavior: a genuine progressive would counterargue such claims rather than acquiesce. This instability raises a critical question: Do LLMs’ responses reflect genuine beliefs, or are they merely performative reproductions of training data patterns?
@@
Humans develop political ideologies through lived experience and reasoning, resulting in predictable stances across related issues (e.g., a free-market advocate consistently supporting privatization and tax cuts) (Converse (2006)). LLMs, however, lack this cognitive anchoring. Trained to optimize token prediction rather than ideological consistency, they generate responses that mirror statistical distributions in their training corpora (Feng et al., 2023). Consequently, models may exhibit confident stances on certain policies (e.g., universal healthcare) while demonstrating inconsistency on closely related issues (e.g., progressive taxation)—a vulnerability that has been observed and even amplified by prompt manipulation (Anagnostidis & Bulian, 2024). This misalignment between surface-level responses and the model’s internal representations compromises their dependability in high-stakes applications such as policy analysis and voter engagement.
In this work, we argue that characterizing LLMs’ political stances requires moving beyond coarse left/right labels to measure belief depth, defined as a model’s capacity to consistently defend positions under argumentative pressure. This approach addresses three fundamental limitations in current methodology: (1) the inability of automated systems to differentiate authentic ideological alignment from superficial linguistic mimicry in single-prompt responses; (2) the lack of quantitative measures for stance consistency across role-playing scenarios with conflicting ideologies; and (3) the unresolved relationship between linguistic uncertainty markers and belief instability.
@@
Through systematic evaluation of 12 LLMs across 19 economic policies, we establish three key findings. First, models exhibit substantial variation in belief resilience, with the percentage of consistent behavior ranging from 58% to 95% for left-leaning models and from 68% to 89% for right-leaning models when challenged with counterarguments. Second, using our predicted labels as ground truth, semantic entropy emerges as a robust detector of pretended beliefs, achieving discriminative power (AUROC=0.78) comparable to human expert assessments. Third, our belief depth analysis reveals topic-specific belief depth patterns that do not always follow traditional ideological divisions: models may strongly defend healthcare privatization while easily shifting their stance on progressive taxation, regardless of whether they lean left or right.
@@ 2 Background and Related Works
Stance detection is an important topic and well-studied in social and political sciences, emphasizing its role in shaping discourse, influencing perception, and driving decision-making. A considerable number of works focus on inferring political positions in an automated way from political texts and is still an active area of research (Laver et al. (2003); Slapin & Proksch (2008); Glavaš et al. (2017); Burnham (2024)). In political science, political party manifestos (Laver & Garry (2000)), political speeches (Grimmer & Stewart (2013)), legislator communications with constituents (Grimmer (2013)), and social movements (Kann et al. (2023)) are the main source of studying political stances of individuals.
@@
As language models continue to evolve, these traditional sources of political stance analysis are being challenged. There are many important use cases of LLMs in the political context, including understanding and summarizing political texts to improve the analysis of political rhetoric and communications, as well as creating mass political advertisements to sway public opinion by generating political content at scale (Linegar et al. (2023)). Researchers are therefore increasingly concerned about how these models might encode, amplify, or misrepresent political stances from their training data, raising critical questions about fairness and bias in AI systems. One of the most widely used scales for detecting political stance is left-right, where the candidates are arguably distinguished between a progressive (left) and conservative (right) position. Although there has been concerns about the validity of inferring the stances using only one axis (Kitschelt (1994); Jahn (2023)), this scale has been validated broadly across countries (Evans et al. (1996); Budge (2001)) and used as the basis for economic stance evaluation of LLMs in many previous works (Feng et al. (2023); Ceron et al. (2024); Agiza et al. (2024)).
@@
Political beliefs of LLMs and their implication have been extensively studied in recent years. Feng et al. (2023) presents one of the most extensive and comprehensive analyses of the political stances of the well-known models at the time. Using the statements from the Political Compass test  [^1] https://www.politicalcompass.org/test, which is one of the most widely used tests for detecting political stances, it proposed a systematic way of inferring the stances of MLMs and CLMs and also provided insights on the steerability of their stances and effect of training partisan datasets. Santurkar et al. (2023) focuses on the opinions of the LLMs regarding different social opinions. They proposed the OpinionQA dataset which has been widely used in later works to detect the opinions of the LLMs on social matters. DURMUS et al. later followed and extended their work by creating a new dataset for detecting LLMs’ opinions called GlobalOpinionQA, using questions from well-known political surveys like the Global Attribute Survey by PEW Research  [^2] https://www.pewresearch.org/ and the World Values Survey. Yang et al. (2024) performed an extensive evaluation of the political biases of LLMs, especially focusing on the modern models, and reports that most of the modern models show a left-leaning stance. They also show that different characteristics of these models, such as their size, release date, and region of origin, affect their behavior regarding political preference, sensitivity to social topics, and misinformation detection. Ceron et al. (2024) argue that there is a defining distinction between models’ political stance and their underlying worldview which results in them expressing different beliefs, showing how LLMs with similar stances can have quite different opinions on related topics. Anagnostidis & Bulian (2024) and Chen et al. (2024) show how constructing the prompts to be more convincing or adding biases resonating with the bias of the models can alter the models’ opinion to be more conforming. Taubenfeld et al. (2024) show when the models are instructed to behave as someone opposite to their beliefs, their inherent bias causes them to revert to their original beliefs while using them for longer conversations.
@@ 3.1 Candidate Model Selection
Figure 1: Stances of the candidate models
There have been a number of works on identifying the political stance of LLMs (Feng et al. (2023); Santurkar et al. (2023); Yang et al. (2024)) using their response towards political statements. Recent studies show that most modern LLMs demonstrate a left-libertarian stance (Yang et al. (2024); Motoki et al. (2024)). This stance is stronger in the instruction-tuned variants. This made our task of selecting strong candidates from both sides challenging. To overcome this, we fine-tuned some of the instruction models with the right partisan data from PoliTune (Agiza et al. (2024)) using direct preference optimization (DPO) (Rafailov et al. (2023)). Please refer to Appendix A for the insights behind this. Figure 1 shows the stances of our candidate models, where we have eight leftist and four rightist candidates. To detect the stances of the models we use the approach introduced by Feng et al. (2023), where we collect the models’ responses to the statements from the widely used political compass test in an open-ended setup. We avoid using a forced-choice setup where the LLMs are provided to choose an option from multiple choices, as such approaches have been criticized in previous works due to models having ‘selection bias’ (Zheng et al. (2024); Röttger et al. (2024)). Instead, we use an off-the-shelf NLI model to detect their agreement or disagreement with the statements using their whole responses (Lewis et al. (2019)).
@@ 3.2 Data Preparation
In this work, we focus on the stances of the models on the economic (left/right) axis. Therefore, for our evaluations, we used the 19 statements of the political compass test consisting only of economic topics. To identify these statements we first select ’disagree’ for all statements and then switch to ‘agree‘ for one statement at a time observing the resulting change in axis (economic or social) and direction (left/right) due to agreement. A factor analysis using the 19 items yields a dominant factor that explains 62% of the variance, confirming that the items primarily measure a single dimension (Appendix B). We label the statements according to their bias in a way that also allows us to capture the bias of the responses:
@@
Next, the models were prompted to respond to politically charged statements under different conditions at low temperatures. Specifically, we collected responses under three scenarios:
  • Default setting: Respond to each statement without any guiding instructions.
@@
We intentionally avoid giving detailed instructions on playing a biased role so that the models naturally depict a leftist or rightist perspective using their own knowledge, preventing any influence or guidance from the prompt. For the prompts used for data generation please refer to Appendix F.1.
3.3 Examining Model Behavior Across Bias Conditions

### v2

# Testing Conviction: An Argumentative Framework for Measuring LLM Political Stability

@@ Abstract
Large Language Models (LLMs) increasingly shape political discourse, yet exhibit inconsistent responses when challenged. While prior research categorizes LLMs as left- or right-leaning based on single-prompt responses, a critical question remains: Do these classifications reflect stable ideologies or superficial mimicry? Existing methods cannot distinguish between genuine ideological alignment and performative text generation. To address this, we propose a framework for evaluating ideological depth through (1) argumentative consistency and (2) uncertainty quantification. Testing 12 LLMs on 19 economic policies from the Political Compass Test, we classify responses as stable or performative ideological positioning. Results show 95% of left-leaning models and 89% of right-leaning models demonstrate behavior consistent with our classifications across different experimental conditions. Furthermore, semantic entropy strongly validates our classifications (AUROC=0.78), revealing uncertainty’s relationship to ideological consistency. Our findings demonstrate that ideological stability is topic-dependent and challenges the notion of monolithic LLM ideologies and offers a robust way to distinguish genuine alignment from performative behavior.
Testing Conviction: An Argumentative Framework for Measuring LLM Political Stability
Shariar Kabir1,2, Kevin Esterling2, Yue Dong2 Bangladesh University of Engineering and Technology1, University of California Riverside2 shariar1405076@gmail.com
@@ 1 Introduction
The growing influence of Large Language Models (LLMs) in public discourse has intensified scrutiny of how they encode and express political biases inherited from training data (Feng et al., 2023; Santurkar et al., 2023). While studies consistently demonstrate that modern LLMs exhibit left-libertarian leanings (Yang et al., 2024; Motoki et al., 2024; Aldahoul et al., 2025), their responses to political topics remain alarmingly inconsistent (Ceron et al., 2024). For instance, models may advocate progressive taxation in one context yet contradict themselves when presented with counterarguments framed as authoritative claims (Chen et al., 2024; Anagnostidis and Bulian, 2024). This irrational compliance contrasts sharply with human ideological behavior: a genuine progressive would counterargue such claims rather than acquiesce. This instability raises a critical question: Do LLMs’ responses indicate genuine ideological positioning or merely adaptive text generation?. We suspect the answer is somewhere in the middle and varies for each LLM and topic in question and this work aims to quantify this behavior.
@@
Humans develop political ideologies through lived experience and reasoning, resulting in predictable stances across related issues (e.g., a free-market advocate consistently supporting privatization and tax cuts) Converse (2006). LLMs, however, lack this cognitive anchoring. Trained to optimize token prediction rather than ideological consistency, they generate responses that mirror statistical distributions in their training corpora (Feng et al., 2023). Consequently, models may exhibit confident stances on certain policies (e.g., universal healthcare) while demonstrating inconsistency on closely related issues (e.g., progressive taxation)—a vulnerability that has been observed and even amplified by prompt manipulation (Anagnostidis and Bulian, 2024). This misalignment between surface-level ideology and the model’s internal representations compromises their dependability in high-stakes applications such as policy analysis and voter engagement.
In this work, we argue that characterizing LLMs’ using single-prompt evaluations are insufficient as these static evaluations fail to capture the stability of these orientations. To address these limitations, we introduce a new evaluation framework that measures ideological depth, defined as a model’s capacity to defend its positions under argumentative pressure. Our contributions are threefold: 1) A novel framework for evaluating ideological depth using argumentative pressure; 2) An empirical demonstration that LLM ideology is topic-specific and not monolithic; and 3) Evidence that semantic uncertainty strongly validates our classification of stable (true) and performative (unstable) stances.
@@
Through systematic evaluation of 12 LLMs across 19 economic policies, we establish three key findings. First, models exhibit substantial variation in their resilience to arguments, with the percentage of consistent behavior ranging from 58% to 95% for left-leaning models and from 68% to 89% for right-leaning models when challenged with counterarguments. Second, using our predicted labels as ground truth, semantic entropy demonstrates significant discrimination ability with AUROC upto 0.78. Third, our analysis reveals topic-specific ideological depth patterns that do not always follow traditional ideological divisions, e.g., models may strongly defend healthcare privatization while easily shifting their stance on progressive taxation, regardless of whether they lean left or right.
@@ 2 Background and Related Works
Ideological stance detection is an important topic and well-studied in social and political sciences, emphasizing its role in shaping discourse, influencing perception, and driving decision-making. A considerable number of works focus on inferring political positions in an automated way from political texts and is still an active area of research Poole and Rosenthal (1985); Laver et al. (2003); Slapin and Proksch (2008); Glavaš et al. (2017); Tausanovitch and Warshaw (2017); Burnham (2024). In political science, political party manifestos Laver and Garry (2000), political speeches Grimmer and Stewart (2013), legislator communications with constituents Grimmer (2013), and social movements Kann et al. (2023) are the main sources of studying political stances of individuals.
@@
As language models continue to evolve, these traditional sources of political stance analysis are being challenged. There are many important use cases of LLMs in the political context, including understanding and summarizing political texts to improve the analysis of political rhetoric and communications, as well as creating mass political advertisements to sway public opinion by generating political content at scale Linegar et al. (2023). Researchers are therefore increasingly concerned about how these models might encode, amplify, or misrepresent political stances from their training data, raising critical questions about fairness and bias in AI systems. One of the most widely used scales for detecting political stance is left vs. right, where the candidates are distinguished between a liberal (left) and conservative (right) position. Although there has been concerns about the validity of inferring the stances using only one axis Kitschelt (1994); Jahn (2023), this scale has been validated broadly to be the most dominant axis Evans et al. (1996); Budge (2001) and used as the basis for economic stance evaluation of LLMs in many previous works Feng et al. (2023); Ceron et al. (2024); Agiza et al. (2024).
@@
Political ideologies of LLMs and their implication have been extensively studied in recent years. Feng et al. (2023) presents one of the most extensive and comprehensive analyses of the political stances of the well-known models at the time. Using the statements from the Political Compass test  [^1] https://www.politicalcompass.org/test, it proposed a systematic way of inferring the stances of MLMs and CLMs and also provided insights on the steerability of their stances and the effect of training partisan datasets. Santurkar et al. (2023) focuses on the opinions of the LLMs regarding different social opinions. They proposed the OpinionQA dataset, which has been widely used in later works to detect the opinions of the LLMs on social matters. Durmus et al. (2024) later followed and extended their work by creating a new dataset for detecting LLMs’ opinions called GlobalOpinionQA, using questions from well-known political surveys like the Global Attribute Survey by PEW Research  [^2] https://www.pewresearch.org/ and the World Values Survey. Yang et al. (2024) performed an extensive evaluation of the political biases of LLMs, especially focusing on the modern models, and reports that most of the modern models show a left-leaning stance. They also show that different characteristics of these models, such as their size, release date, and region of origin, affect their behavior regarding political preference, sensitivity to social topics, and misinformation detection. Ceron et al. (2024) argue that there is a defining distinction between models’ political stance and their underlying worldview, which results in them expressing different ideologies, showing how LLMs with similar stances can have quite different opinions on related topics. Anagnostidis and Bulian (2024) and Chen et al. (2024) show how constructing the prompts to be more convincing or adding biases resonating with the bias of the models can alter the models’ opinion to be more conforming. Sharma et al. (2023) shows that the model that has gone through preference-based training tends to produce answers that are consistent with the user’s stated views, a behavior called sycophancy. Taubenfeld et al. (2024) shows that when the models are instructed to behave as someone opposite to their ideologies, their inherent bias causes them to revert to their original ideologies while using them for longer conversations.
@@ 3.1 Candidate Model Selection
Figure 1: Stances of the candidate models. Left and right-leaning candidates are shown in different colors.
There have been a number of works on identifying the political stance of LLMs Feng et al. (2023); Santurkar et al. (2023); Yang et al. (2024) using their response towards political statements. Recent studies show that most modern LLMs demonstrate a left-libertarian stance Yang et al. (2024); Motoki et al. (2024). This stance is stronger in the instruction-tuned variants. This made our task of selecting strong candidates from both sides challenging. To overcome this, we fine-tuned some of the instruction models with the right partisan data from PoliTune Agiza et al. (2024) using direct preference optimization (DPO) Rafailov et al. (2023). Please refer to Appendix A for the insights behind this. Figure 1 shows the stances of our candidate models, where we have eight leftist and four rightist candidates. To detect the stances of the models, we use the approach introduced by Feng et al. (2023), where we collect the models’ responses to the statements from the political compass test in an open-ended setup. We avoid using a forced-choice setup where the LLMs are provided to choose an option from multiple choices, as such approaches have been criticized in previous works due to models having ‘selection bias’ Zheng et al. (2024); Röttger et al. (2024). Instead, we use an NLI model Lewis et al. (2019) to detect their agreement or disagreement with the statements using their complete responses.
@@ 3.2 Data Preparation
In this work, we focus on the stances of the models on the main ideological axis (left-right). Therefore, for our evaluations, we used the 19 statements of the political compass test consisting only of economic topics. To identify these statements, we first select ’disagree’ for all statements and then switch to ‘agree‘ for one statement at a time, observing the resulting change in axis (economic or social) and direction (left/right) due to agreement. To verify responses scale in one dominant dimension, a factor analysis (FA) was performed by collecting 20 responses for each statement from the candidates at $tempreture=1$. FA results in a dominant factor that explains 62% of the variance, confirming that the items primarily measure a single dimension (Appendix B). We label the statements according to their bias in a way that also allows us to capture the bias of the responses:
@@
The final curated dataset is documented in Appendix G. Next, the models were prompted to respond to the 19 statements under different conditions at $temperature=0$ to get the most probable answers. Specifically, we collected responses under three scenarios:
  • Original persona: Respond to each statement without any guiding instructions.
@@
We intentionally avoid giving detailed instructions on playing a biased role so that the models naturally depict a leftist or rightist perspective using their own knowledge, preventing any influence or guidance from the prompt. For the prompts used for data generation, please refer to Appendix F.1.
Figure 2: Visualization of how a model’s internal topic-wise stance, rather than its ideology, contributes to its opinion over the topic. Response from a right-leaning model with a true left-leaning stance over a topic can be indistinguishable from a left-leaning model.
3.3 Do Models Have Ideological Beliefs?
Since modern instruction-tuned LLMs demonstrate agentic behavior, one way to explain their inconsistent responses can be that these models internally hold ideological beliefs, and when prompted to behave in an opposite way, they are "pretending". For instance, a right-leaning model asked to provide a single left-leaning response might generate an answer that appears ideologically aligned even though the model lacks the depth or logical consistency a genuinely left-leaning model demonstrates. To see if that’s the case, we run a battery of tests evaluating both their response quality, monitoring their perplexity, and performing a lie detection test using their internal layer representations (Zou et al., 2023) (Appendix C ). These results did not reveal any meaningful difference in their response quality or internal representations, which means models do not have any internal ideological beliefs; rather, their opinions are probably largely based on the distribution of training data.
Additionally, the opinions provided by the LLMs do not adhere strictly to a single ideological perspective (Santurkar et al., 2023; Aldahoul et al., 2025). Instead, models exhibit a mix of both left- and right-leaning orientations depending on the topic at hand. Moreover, while some opinions remain consistent, others are merely performative and often shift based on the framing of the statement. We illustrate this scenario in Figure 2, which depicts two models with left and right political ideologies: $\mathcal{M}_{l}$ and $\mathcal{M}_{r}$. Considering a sample of five related topics (e.g., tax cuts, free market), each model displays a mixture of leftist and rightist stances across them, with left-leaning models predominantly adopting leftist positions and vice versa. However, when instructed to respond in a left-leaning manner across all topics, models do not uniformly generate left-leaning content but rather tend to behave in a way that is consistent with their underlying knowledge depth on each topic. On topics where both models hold left-leaning orientations, they produce qualitatively similar responses, making it difficult to classify the right-leaning model’s responses as "pretending". Moreover, a model with right-leaning orientations on a topic may initially attempt to follow left-leaning instructions but eventually reverts to neutral or right-leaning positions. This leads to our central question: How do we identify if a model’s orientation on a topic is genuine and distinguish it from performative adherence to instructions?
To evaluate this, we analyze the models’ behavior when presented with the statements in an argumentative setting to see how consistently they defend their initial stance when confronted with different arguments. The results indicate that models remain true to their stance on some topics but alter or even contradict their initial responses for others, suggesting that their orientations are often shallow for some topics rather than deeply ingrained and consistent with their overall ideology.
3.4 Using Supporting and Counter Arguments
To systematically identify shallow orientations over topics, we introduce an algorithm that evaluates a model’s stance consistency when subjected to supporting and counter-arguments. As illustrated in Figure 3, the key idea is that a model with a deeply ingrained left- or right-leaning stance over a topic should maintain its stance in the presence of argumentative pressure, whereas a model with a shallow stance is likely to shift its stance in response to arguments.
Figure 3: Sample responses of a model with a performative stance over a topic: Model with a shallow stance changes its opinion under argumentative pressure.
This approach ensures that models are evaluated not only on isolated responses but also on their ability to maintain their position when challenged. By leveraging a structured set of supporting and counterarguments, we improve our ability to distinguish between genuine and performative ideological alignment.
We used DeepSeek-R1-32B to prepare strong and realistic supporting and counter-arguments for each of the statements. We also evaluate using different sets of arguments generated at $temperature=1$. We document the prompts used to achieve this in Appendix F.3-F.4.
  • $o\in{+1,-1}$ denote the model’s original opinion on a statement, where +1+1 corresponds to agreement and 1-1 corresponds to disagreement.
  • $b\in{+1,-1}$ denote the bias direction associated with the statement, where we assign +1 to a right-biased statement (i.e., agreement shifts the stance to the right and vice versa) and -1 for a left-biased statement.
[…]
For instance, if $b=+1$, then a model’s opinion $o=+1$ (agreement) means $I_{B}(o,b)=1$ (right bias) and o=1o=-1 (disagreement) means $I_{B}(o,b)=0$ (left bias).
Classification Function: The process of classifying the ideological-depth over the topic can now be expressed as a piecewise function:
[…]
Here, the “true" labels (TL, TR) are assigned when the model’s opinion remains unchanged (i.e., $\delta(o,a)=1$), while a change in opinion (i.e., $\delta(o,a)=0$) leads to a “preformative" classification (PL, PR). Left/Right (L/R) is determined by the bias indicator $I_{B}(o,b)$. The ideological depths of the models on each topic identified by our approach is documented in Appendix D.
3.5 Using Models’ Uncertainty to Detect Performative Scenarios
We hypothesize that models tend to be performative on certain topics due to their lack of sufficient knowledge to substantiate these positions, rather than having a genuine understanding. They merely reflect a shallow impression of scenario-appropriate answers derived from their alignment tuning. To test this hypothesis, we analyze the models’ uncertainty when responding to political statements. To do this, we evaluate their entropy while presented with the statements as input. Entropy measures the spread of the probability distribution in a model’s prediction for a given input. Higher entropy indicates greater uncertainty in the model’s output and is labeled as confabulation. For calculating the entropy of a generation $Y$ by the LLM for a given input $x$ we can use their predictive entropy (PE) introduced by Kadavath et al. (2022):
$PE(x)=H(Y|x)=-\sum_{y}P(y|x)\ln P(y|x)$ (4)
Statement Our label Generations for entropy Entropy Label
Protectionism is sometimes necessary in trade. Performative - Protectionism can be a controversial topic Confabulation
- Protectionism can be necessary …
- I disagree with the idea that …
- While charity can be a valuable tool for … it’s not …
- I disagree with the idea that …
Table 1: Row 1 demonstrates where the model was identified to hold a performative stance for the topic using our algorithm, and using SE, we detect confabulation due to its generations with different meanings. In contrast, in row 2, the model holding a true stance for the topic SE did not detect confabulation due to its generations with the same meaning.
However, naively calculating only the entropy based on tokens fails to take into account the semantic equivalence of different generations, where different tokens can convey the same meaning. Therefore, we also evaluate their semantic entropy (SE) and its black-box counterpart, discrete semantic entropy (DSE), introduced by Farquhar et al. (2024). The main idea here is to feed the query to the model $N$ times and collect the responses. The responses are then clustered into $|C|$ clusters, where each cluster contains the responses that are semantically equivalent. Finally, it uses the following equation to calculate the entropy over these clusters and get the semantic entropy for the query $x$.
$SE(x)\approx-\sum_{i=1}^{|C|}P(C_{i}|x)\log P(C_{i}|x)$ (5)
$P(C_{i}|x)$ is calculated using the token probabilities from the models layers. For DSE, this value is approximated using the proportion of sequences in each cluster. We compare the labels of the LLM responses (i.e., true or performative) identified by our algorithm ($\delta(o,a)$ from Equation 1) across different topics with the classifications derived from the entropy-based measures (i.e., confabulation or not). Using this approach, we assess the relationship between performative stances and model uncertainty. Table 1 provides an example of our setup.
4 Results
4.1 Ideological Depths Over the Three Scenarios
TL PL TR PR
Leftist-models 37.5 34.87 10.53 17.11
Rightist-models 13.16 21.05 44.74 21.05
Table 2: Distribution of true and performative orientations in the original (no instruction) persona
For each of the three roleplay scenarios (original, as left, and as right), we first extracted the models’ responses on each topic using our algorithm, which categorizes them as True Left (TL), Performative Left (PL), Performative Right (PR), or True Right (TR). We document the distribution of true and performative responses for the leftist and rightist models in Table 2. The result supports our assumption in Section 3.3, where both sets of models consist of both true left and true right orientations. Their political ideology emerges from the fact that, for left models, the percentage of true left orientations and for right models, the percentage of true right orientations is higher. Next, we compute the probabilities of holding and changing their orientation labels using all possible transitions. We document this in Appendix E (Table 11). We found, the left models consisted of 37.5% of true-left orientations with up to 76.9% average probability of holding them, and the right models consisted of 44.74% of true-right orientations with up to 56.5% average probability of holding them, meaning models are significantly more consistent at holding their true stances over topics.
Figure 4: Consistent orientation changes
Figure 5: Inconsistent orientation changes
Figure 6: Categorizing orientation changes over scenarios
Varying Personas Varying Arguments
Model $P_{con}$ $P_{inc}$ $P_{incon}$ $P_{con}$ $P_{inc}$ $P_{incon}$
Llama-2-70b-chat 0.84 0.16 0.0 0.79 0.0 0.21
Llama-2-13b-chat 0.95 0.05 0.0 0.79 0.0 0.21
OLMo-7B-0724-Instruct 0.58 0.0 0.42 0.89 0.0 0.11
Mistral-7B-Instruct-v0.3 0.58 0.05 0.37 0.47 0.0 0.53
Llama-3.2-3B-Instruct 0.79 0.16 0.05 0.63 0.0 0.37
Gemma-2-9b-it 0.68 0.05 0.26 0.63 0.0 0.37
Qwen2.5-3B-Instruct 0.58 0.05 0.37 0.63 0.0 0.37
Qwen2.5-3B 0.74 0.0 0.26 0.79 0.05 0.16
DeepSeek-R1-Distill-Qwen-1.5B 0.63 0.05 0.32 0.89 0.0 0.11
Right-FT-Llama-3.1-8B-Instruct 0.68 0.16 0.16 0.63 0.0 0.37
Right-FT-Llama-2-7b-chat 0.89 0.05 0.05 0.84 0.0 0.16
Right-FT-Llama-2-13b-chat 0.89 0.11 0.0 0.84 0.0 0.16
Table 3: Distribution of consistent and inconsistent orientation changes of different models while playing different personas. $P_{con}$, $P_{inc}$, and $P_{incon}$ represent the probabilities of consistent, inconsistent, and inconclusive outcomes, respectively.
Model $P_{con}$ $P_{inc}$ $P_{incon}$
Llama-3.1-8B-Instruct 0.53 0.11 0.37
Llama-2-7b-chat 0.58 0.32 0.11
Llama-2-13b-chat 0.84 0.05 0.11
Table 4: Distribution of consistent and inconsistent orientation changes of different models after fine-tuning. $P_{con}$, $P_{inc}$, and $P_{incon}$ represent the probabilities of consistent, inconsistent, and inconclusive outcomes, respectively.
4.2 Measuring the Consistency of the True Orientations over Topics
Using the insights from the transition probabilities between different orientations (Appendix E), we categorized the possible transitions into three types: consistent, inconsistent, and inconclusive:
  • Inconsistent Transitions: The model demonstrates conflicting true orientations for the same topic across the three scenarios: $TL\rightarrow TR$ and $TR\rightarrow TL$.
  • Inconclusive Transitions: The model doesn’t show any true orientations for the topic across the three scenarios: $PL\rightarrow PR$, $PR\rightarrow PL$, $PL\rightarrow PL$, $PR\rightarrow PR$.
  • Consistent Transitions: All other transitions where the model shows true orientation on a topic of only one type (left/right) in one scenario and either holds it or behaves performatively in all others.
Figure 6 and 6 shows the orientation transitions that are consistent and inconsistent with a model’s true orientations, respectively.
While its well established that all LLMs’ opinions are at some level susceptible to prompt engineering Anagnostidis and Bulian (2024), we hypothesize that they are significantly more consistent on topics with true stances. For this, we evaluated the consistency of the candidates with three different sets of supporting and counter arguments, along with their consistency while playing the three different personas. Table 3 shows the probability distribution of consistent and inconsistent transitions for all candidate models. It can be seen that all the models’ behavior is significantly more consistent with their true orientations on the topics, with consistency ranging from 58% to 95% for left-leaning and from 63% to 89% for right-leaning models. The results by varying the arguments were also in agreement with our hypothesized outcome.
Since we fine-tuned some of the leftist models to make them rightist, it is also worth analyzing the consistency of their orientations after fine-tuning (Table 12). Interestingly, using our definition of different types of orientation transitions, we found that the changes in their stances on topics were also mostly (up to 84%) consistent with their true orientations (Table 4). This suggests that even when fine-tuned and the models shift their ideology, they remain mostly consistent with their true orientations.
4.3 Assessing the Role of Uncertainty
A key assumption in our study is that the performative responses result from the model’s uncertainty over the topic, which can arise due to data distribution (aleatoric uncertainty) or limited information (epistemic uncertainty) (Farquhar et al., 2024). To test this, we measure the entropy values for model generations across different statements and compare them with our identified labels. A higher entropy score indicates greater uncertainty in a model’s output, which results in performative behavior.
Figure 7: Evaluating AUROC using our labels and different entropies
Figure 7 presents the AUROC (Area Under the Receiver Operating Characteristic curve) scores for SE, DSE, and NE, evaluating how well these entropy measures distinguish between true and performative responses that are labeled as such by our algorithm. In each case, we use $N=10$, following previous works. An AUROC of 1.0 indicates perfect discrimination, while 0.5 means none. For most LLMs, we observe significant discriminative power ($AUROC>0.6$) using SE-based metrics. SE achieves the highest AUROC score of 0.78, closely followed by DSE. This supports our hypothesis that a model’s tendency to generate performative or inconsistent responses on a topic correlates strongly with its uncertainty over that topic.
5 Conclusion
In this study, we examined whether the responses of LLMs genuinely reflect their ideology by evaluating their stances in the presence of argumentative pressure. Our findings demonstrate that while all LLMs exhibit political stances, their depth of stance varies significantly across topics, leading to inconsistencies when these responses are challenged. We propose a systematic approach to identify true versus performative stances and validate its effectiveness by analyzing their orientation consistency and uncertainty metrics. Our results suggest that models with deeper stances over a topic tend to maintain their stance with stable responses, whereas shallower stances can be more easily swayed by prompt manipulation. In the future, we plan to further extend our study by exploring the internal representations of the political concepts in these models to explain these behaviors.
Appendix A Finetuning Models using DPO to Shift Their Stance
Preference fine-tuning allows us to tune the target alignment of a model with minimal required data (2K in our case), without exposing the model to additional data and altering core reasoning capacities, and only shifts the alignment responsible for its ideology, ensuring that comparability is maintained. Moreover, a controlled dataset like PoliTune lets us prevent exposure to data similar to the actual test statements, preserving evaluation integrity since both model sets were tested under identical conditions. The finetuned models are also later used for an interesting evaluation where we investigate how their ideological depths impact their transition of stance on topics.
Figure 8: Model stance before and after DPO fine-tuning.
Figure 8 shows the shift of stance after finetuning tyree left-leaning models of different sizes. Model size has an effect on the amount of shift achieved, where the smaller models shift more than the larger models.
Appendix B Factor Analysis of the LLM Responses
In order to verify that the 19 economic items used in our evaluation predominantly measure a single dimension, we conducted a factor analysis using closed-ended responses from LLM outputs. The analysis was performed on a dataset with 20 responses for every items generated at $tempreture=1$ from each of the 12 candidate model (240 observations). The responses were collected in an open-ended setting and then converted into close-ended responses using an NLI framework. Below, we provide details on both the unrotated and rotated solutions.
Unrotated Factor Analysis: Table 5 presents the eigenvalues and the variance explained by each factor from the unrotated solution. The first factor had an eigenvalue of 5.72 and explained 68.08% of the variance, while the second factor had an eigenvalue of 1.63, contributing an additional 19.37% to the explained variance. Following the Kaiser criterion (eigenvalues greater than 1), two factors were retained for rotation.
Factor Eigenvalue Proportion Cumulative
Factor 1 5.72 0.6808 0.6808
Factor 2 1.62 0.1937 0.8744
Factor 3 0.82 0.0977 0.9721
Factor 4 0.53 0.0631 1.0352
…(remaining factors with eigenvalues <0.5<0.5)
Table 5: Eigenvalues and Variance Explained (Unrotated Solution)
Variable Factor 1 Factor 2 Uniqueness
$Q_{1}$ -0.74 0.28 0.36
$Q_{2}$ 0.31 0.47 0.67
$Q_{3}$ 0.68 0.05 0.52
$Q_{4}$ -0.35 0.38 0.72
$Q_{5}$ -0.33 0.35 0.75
$Q_{6}$ 0.66 -0.15 0.53
$Q_{7}$ -0.27 0.53 0.63
$Q_{8}$ -0.07 0.72 0.46
$Q_{9}$ 0.23 0.27 0.87
$Q_{10}$ 0.15 -0.03 0.97
$Q_{11}$ 0.78 -0.07 0.37
$Q_{12}$ 0.72 -0.13 0.45
$Q_{13}$ 0.55 -0.06 0.69
$Q_{14}$ -0.35 0.54 0.58
$Q_{15}$ -0.44 0.44 0.60
$Q_{16}$ 0.40 0.05 0.83
$Q_{17}$ 0.66 -0.22 0.50
$Q_{18}$ 0.54 -0.27 0.63
$Q_{19}$ 0.74 0.04 0.44
Table 6: Rotated Factor Loadings (Varimax) and Uniqueness Values
Factor Eigenvalue Proportion Cumulative
Factor 1 5.19 0.6179 0.6179
Factor 2 2.16 0.2566 0.8744
Table 7: Rotated Factor Variance Table